# Training Language Models to Explain Their Own Computations

### v2

@@ 5.2 Input Ablations: Training LMs to Describe their Decision Rules
Similar to the activation patching experiments, we train two models—Llama-3.1-8B-Instruct and Qwen3-8B—on their own behaviors and the behaviors of each other as the target models. Using the methods described in Section˜2.4, we train models to explain how withholding or including parts of target models’ input would affect the output prediction.
@@
Following recent work used to study faithful chain-of-thought (chen2025reasoningmodelsdontsay), we give $\mathcal{M}$ inputs in the form of a multiple-choice question $c$ with an injected hint $\tilde{x}$, i.e. $x=[c,\tilde{x}$], and train the explainer to predict whether $\mathcal{M}$ ’s most likely next token will change to $\tilde{x}$. Previously, chen2025reasoningmodelsdontsay found that models fail to verbalize $\tilde{x}$ in their chain-of-thought, despite $\tilde{x}$ being critical to its prediction ($\mathcal{M}([c,\tilde{x}])\neq\mathcal{M}(c)$). Thus, we investigate whether models can be trained to detect when and how their prediction will change depending on $\tilde{x}$’s presence.
We use MMLU (hendryckstest2021) as the source of multiple-choice reasoning question, and inject hints of the form $\tilde{x}$=Hint: A” to the end of the question. We randomly select $\tilde{x}$ as one of the four multiple-choice options. To make the has-changed prediction nontrivial, we construct the hint prompt such that we have a roughly equal number of samples that change and do not change according to the hint. See Section˜E.3 for details.
Metrics.
Recall from Section˜2.4 that the explainer’s prediction involves two parts, similar to activation patching: whether the target’s prediction will change when the hint is removed (e.g. output changes / output remains unchanged), and the new prediction / content of the change (e.g. to (A)). Thus, we use the same three metrics as in activation patching (§5.1).
Baselines.
Similar to Section˜5.1, we include untrained explainer baselines. The prompt for then untrained models can be found in Section˜F.3.
Target Explainer Exact Match Has-Changed F1 Content Match
Llama-Inst. Llama-Inst. 18.2±1.0 58.2±1.3 24.9±1.2
Qwen3-8B 14.6±0.9 52.4±1.2 22.9±1.1
Llama-Inst. Llama-Inst.
21.0±1.1 45.5±1.3 36.7±1.3
Qwen3-8B Qwen3-8B 62.7±1.3 73.4±1.1 70.5±1.2
Llama-Inst. 33.1±1.3 59.4±1.3 48.3±1.3
Qwen3-8B Qwen3-8B
30.2±1.2 53.2±0.5 39.9±1.3
Table 3: An explainer model can best explain its own input ablation outcomes. We train Qwen3-8B and Llama-3.1-8B to verbalize their own input ablation outcomes, and compare against training to verbalize each other’s outcomes. We also report untrained models explaining themselves. Note this table is sorted by explainer rather than target unlike previous tables. Scores (mean $\pm$ standard error) are reported for three metrics—exact match, has-changed F1, and content prediction. For each explainer model and metric, we bold all entries not significantly different from the best entry (paired $t$-test, $p\geq 0.05$). We find evidence of a different version of privileged access: models are better at explaining themselves than other models.
@@ Results
are shown in Table˜3. We find a different pattern of privileged access in this case, as Llama-3.1-8B appears to be an overall a much weaker explainer of both itself and of Qwen. However, when the explainer (instead of target) is held constant, both models are better at explaining themselves than explaining others: Llama is better at explaining Llama than Qwen (despite being a worse explainer overall), and vice versa. One way of viewing this is that privileged access is a significant first-order term in explainer performance, but can in some cases be overcome by other factors, such as a significant gap in the task-relevant capabilities of the explainer model.
One indication of Llama’s lack of capability comes from comparison to its untrained version: while Llama can learn to better predict its own has-changed value, its content match remains around random (25%) even after training. [^12] Although untrained Llama’s content match is above random, a closer investigation reveals that it’s simply predicting the same content as the hint 81.9% of the time. This results in an above-random content match because if the model prediction remains unchanged, it typically is because the model’s original prediction matches the hint; we do not take the metric as evidence that untrained Llama is doing something non-trivial. If we only look at the has-changed F1, Llama’s ability to explain itself is on par with Qwen’s ability to explain it, despite Llama being a weaker overall explainer.
Finally, we examine untrained Qwen’s predictions and find that it a has-changed value of True only 0.4% of the time, and most of its content match comes from correctly predicting the unchanged answers (91.0% content match) compared to its changed answers (14.3% content match). This is consistent with findings on reasoning models in prior work: chen2025reasoningmodelsdontsay finds that no pre-trained LM would ever verbalize that it utilizes the hint, even when it did, similar to our finding that untrained Qwen never answers that its response changes based on the hint. This indicates that explicit fine-tuning is essential to elicit models to verbalize their true decision process.
Taken together, results on these two tasks show that:
Privileged access extends across tasks, but requires a certain model capacity; some models may be (model-agnostic) poor explainers of certain types of computation.
@@ 6.1 Chain-of-Thought Faithfulness
Language models can be asked to verbalize their thought processes through chain-of-thought prompting, which offers a way for external monitoring (korbak2025chainthoughtmonitorabilitynew; baker2025monitoringreasoningmodelsmisbehavior), but prior work has found that these verbalizations can be unfaithful to their true decision-making processes (turpin2023language; lanham2023measuringfaithfulnesschainofthoughtreasoning; barez2025cot; chen2025reasoningmodelsdontsay). This line of work directly inspires our hint ablation task setup, while our work attempts to remedy the unfaithfulness through a training objective that enforces consistency between model verbalization and their actual mechanisms.
@@ 6.2 Mechanistic Interpretability Methods
The field of mechanistic interpretability sets out to (1) describe individual features and (2) connect them to construct a causal circuit that performs a certain task. While many researchers have found success doing so manually in a fine-grained way both on the neuron level (gurnee2024universal) and the circuit level for toy tasks (wang2023interpretability; nanda2023progressmeasuresgrokkingmechanistic), these methods have been challenging to scale and generalize to new models and tasks (sharkey2025open). To address scalability, automated feature description pipelines (hernandez2022natural; bills2023language; choi2024automatic; paulo2025automatically) and circuit discovery techniques (conmy2023automated; syed-etal-2024-attribution; hanna2024have; hsu2025efficient) have been introduced that require substantial computational resources. Designing a scalable interpretability method that balances correctness/robustness and efficiency is an active area of research, one of which is attribution patching that use gradients to approximate important components to the output (nanda2023attribution; syed-etal-2024-attribution).
Closely related to the current study, one family of scalable interpretability methods attempts to leverage models’ own verbalization ability without any additional training. LogitLens (nostalgebraist2020interpreting), PatchScopes (ghandeharioun2024patchscopes), SelfIE (chen2024selfie) enable models to self-interpret zero-shot, but such methods require significant hyperparameter tuning (kharlapenko2024self).  pan2024latentqa introduces Latent Interpretation Tuning, which fine-tunes models to answer questions about other models’ inputs via their hidden representations. We expand on this line of work by training models to self-verbalize their internals based on data collected from existing interpretability techniques that provide finer-grained pictures of models’ internal computations.
Finally, explanations based on special interpretability techniques are difficult for non-expert users to access, potentially requiring developers integrate many specialized interfaces (viégas2023modelusermodelexploring; chen2024designingdashboardtransparencycontrol). Our work offers an alternative path toward user understanding, in which explanations can be provided “in-band” rather than via an external interface.
@@ 6.3 Introspection & Metacognition
Recent work investigates whether models possess metacognition or introspective abilities (binder2025looking; treutlein2024connecting; laine2024memyselfaisituational; comsa2025doesmakesensespeak; plunkett2025selfinterpretabilityllmscomplexinternal). A central debate in this literature concerns whether models have privileged access to their internals, or whether introspection merely reflects their strong predictive capacity to learn external correlations (song2025language; song2025privilegedselfaccessmattersintrospection; li2025do). We add to the literature by assessing models’ introspective abilities on their own mechanisms derived from mechanistic interpretability methods.
7 Discussion: Self-Consistency as a Framework for Interpretability and Faithfulness
@@
This paper explores improving model faithfulness and interpretability by training models to generate explanations consistent with outcomes of interpretability procedures. Across two question types, we find evidence that generated explanations take advantage of models’ privileged access to their own internal mechanisms, and on the third, we find evidence of explainer models having privileged target access to their own internals. Quantifying how factors such as explainer model capacity, target model complexity, task difficulty, etc. influence the emergence or manifestation of privileged access remains an important question for future work.
@@
In a general sense, our results suggest that objectives that enforce self-consistency between models’ behavior and explanations may provide an avenue towards more scalable interpretability methods, faithful explanations, and alignment guarantees. The experiments in this paper derived (self-)explanation objectives from feature description, activation patching and input ablation techniques. Generalized versions of these objectives could in principle be applied to any method for generating ground-truth descriptions of model computation or behavior, e.g. to describe inferred user traits, surface adversarial triggers, or summarize influential training examples (chen2024designingdashboardtransparencycontrol; Wallace2019Triggers; pmlr-v70-koh17a).
One natural extension would also optimize for consistency in a stronger sense: rather than having fine-tuned explainer models explain static target models, enforce that the target model is always identical to the explainer model. Perhaps more interestingly, these objectives could be applied in the reverse direction, fine-tuning models so that their behavior matches generated explanations, which might be normatively correct even if inaccurate as descriptions of models themselves. Applied in this way, consistency fine-tuning could provide a new mechanism for “unsupervised alignment” without relying on large preference datasets or reinforcement learning procedures.
Acknowledgments
We would like to thank Dami Choi for technical support at early stages of this project. This work was supported by the Open Philanthropy Foundation and the MIT Quest for Intelligence. BZL is additionally funded by a Clare Boothe Luce Fellowship, and JA is funded by a Sloan Fellowship.
Appendix A Feature Descriptions: Additional Experimental Details
A.1 Explainer models
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
A.2 Metric Details: LM Judge
LM Judge Prompt.
@@ A.2 Metric Details: LM Judge
Appendix B Feature Descriptions: Additional Results
B.1 Gemma-2-9B as Target Model
We replicate a subset of the evaluations in Table˜1 using Gemma-2-9B as the target model. We select Gemma-2-9B because its SAE features are also publicly available via GemmaScope (lieberum-etal-2024-gemma), and downloadable via Neuronpedia (neuronpedia). Results can be found in Table˜4.
@@ B.1 Gemma-2-9B as Target Model
B.2 Quantifying privileged access via activation alignment
@@
For each similarity metric and each explainer model (fixing 3.1-8B as the target model), we provide raw similarity values in Table˜5. We then plot explainer similarities vs. explainer performance for each of the four (metric, feature type) pairings, as described in Section˜3.2. Plots can be found in Figure˜6. We find that explainer performance tracks the activation similarity of the explainer, indicating alignment between activations predicts verbalization capability.
@@
B.3 Additional Scaling Curves
We plot scaling curves for Llama-3.1-8B, Qwen3-8B, and nearest neighbors as explainer models for the Llama-3.1-8B target model. After training on various amounts of training data, we plot the simulator scores for each of the three feature types described in Section˜3.2. These plots can be found in Figure˜7.
[…]
We find that the general trend of self-explanation (Llama-3.1-8B) being significantly more data efficient than explaining with another model (Qwen-3-8B) or using nearest neighbors holds on all three types of features when scored using simulator correlations, similar to our findings using LM judge on SAE features in Section˜4.
B.4 Is Layer Annotation Necessary to Describe SAE Features?
@@
B.5 Error Analysis
We perform a human qualitative analysis on 100 test set outputs from the explainer model that received a score of 0 from the LM judge in Table˜1. Our analysis reveals that sources of error extend beyond the explainer model itself and can be attributed to multiple stages of the evaluation pipeline, including (1) the original feature labels generated by the LM based on activation exemplars and (2) the validity of the LM judge’s assessment when comparing the explainers’ output against the original labels. Errors from these sources do not reflect failures of the explainer model.
Appendix C Activation Patching: Additional Results
C.1 Ablation Analysis
We analyze why ablating the layer and token in Table˜2 makes very little difference in the final result. We train Llama-3.1-8B to decode the layer and token information from the input context and activation value alone, i.e. given the prompt
@@ C.1 Ablation Analysis
Appendix D Explainer Model Prompts

### v3

@@ 5.2 Input Ablations: Training LMs to Describe their Decision Rules
Using the methods described in Section˜2.4, we train models to explain how withholding or including parts of target models’ input would affect the output prediction.
@@
Following recent work used to study faithful chain-of-thought (Chen et al., 2025), we give $\mathcal{M}$ inputs in the form of a multiple-choice question $c$ with an injected hint $\tilde{x}$, i.e. $x=[c,\tilde{x}$], and train the explainer to predict whether $\mathcal{M}$ ’s most likely next token will change to $\tilde{x}$. Previously, Chen et al. (2025) found that models fail to verbalize $\tilde{x}$ in their chain-of-thought, despite $\tilde{x}$ being critical to its prediction ($\mathcal{M}([c,\tilde{x}])\neq\mathcal{M}(c)$). Thus, we investigate whether models can be trained to detect when and how their prediction will change depending on $\tilde{x}$’s presence.
We use MMLU (Hendrycks et al., 2021) as the source of multiple-choice reasoning question, and inject hints of the form $\tilde{x}$=Hint: A” to the end of the question. We randomly select $\tilde{x}$ as one of the four multiple-choice options. To make the has-changed prediction nontrivial, we construct the hint prompt such that we have a roughly equal number of samples that change and do not change according to the hint. See Section˜G.3 for details.
@@ Results
for input ablation are also shown in Table˜2. We find similar evidence for privileged access, where Qwen3-8B is best explained by itself than other models. Finally, we examine untrained Qwen’s predictions and find that it a has-changed value of True only 8.6% of the time, and most of its content match comes from correctly predicting the unchanged answers (96.4% content match) compared to its changed answers (10.3% content match). This is consistent with findings on reasoning models in prior work: Chen et al. (2025) finds that no pre-trained LM would ever report using a hint. This indicates that explicit fine-tuning is essential to elicit faithful explanations of decision rules. Taken together, results on these two tasks show that:
Privileged access holds across tasks.
@@ 6.1 Chain-of-Thought Faithfulness
Language models can be asked to verbalize their thought processes through chain-of-thought, which offers a way for external monitoring (Korbak et al., 2025; Baker et al., 2025), but prior work has found that these verbalizations can be unfaithful to their true decision-making processes (Turpin et al., 2023; Lanham et al., 2023; Barez et al., 2025; Chen et al., 2025). This directly inspired our hint ablation task setup, where our work attempted to remedy unfaithfulness through training directly on the models’ own decision rules.
@@ 6.2 Mechanistic Interpretability Methods
Mechanistic interpretability sets out to (1) describe individual features and (2) connect them to construct a causal circuit that performs a certain task. While many researchers have found success doing so manually (Gurnee et al., 2024; Wang et al., 2023; Nanda et al., 2023), these methods have been challenging to scale and generalize to new models and tasks (Sharkey et al., 2025). To address this, automated feature description pipelines (Hernandez et al., 2022; Bills et al., 2023; Choi et al., 2024; Paulo et al., 2025) and circuit discovery techniques (Conmy et al., 2023; Syed et al., 2024; Hanna et al., 2024; Hsu et al., 2025) have been introduced, though these often require substantial compute. Researchers are continuing to actively develop scalable interpretability methods that balance correctness and efficiency (Nanda, 2023; Syed et al., 2024; Shaham et al., 2025).
Closely related to the current study, LogitLens (nostalgebraist, 2020), PatchScopes (Ghandeharioun et al., 2024), SelfIE (Chen et al., 2024a) enable models to self-interpret zero-shot, but require significant hyperparameter tuning (Kharlapenko et al., 2024). Latent Interpretation Tuning (Pan et al., 2024) fine-tunes models to answer questions about other models’ inputs via their activations, and later works have scaled this method up for various explanation settings (Karvonen et al., 2026; Choi et al., 2025; Huang et al., 2025)Goel et al. (2025) introduces a trained LoRA adaptor to verbalize the weight difference of finetuned models. More recently, We train models to self-verbalize their internal computations using finer-grained data collected from interpretability techniques.
Finally, explanations based on special interpretability techniques are difficult for non-expert users to access, potentially requiring developers create specialized interfaces (Viégas & Wattenberg, 2023; Chen et al., 2024b). Our work offers an alternative path toward useability, in which explanations can be provided “in-band” rather than via an external interface.
@@ 6.3 Introspection & Metacognition
Recent work investigates whether models possess metacognition or introspective abilities (Binder et al., 2025; Treutlein et al., 2024; Laine et al., 2024; Comsa & Shanahan, 2025; Plunkett et al., 2025; Lindsey, 2025). A central debate in this literature concerns whether models have privileged access to their internals, or whether introspection merely reflects their strong predictive capacity to learn external correlations (Song et al., 2025a, b; Li et al., 2025). We assess models’ introspective abilities on their own mechanisms derived from mechanistic interpretability methods.
7 Discussion: Self-Consistency as a Framework for Interpretability
@@
This paper explores improving model faithfulness and interpretability by training models to generate explanations consistent with outcomes of interpretability procedures. Across three question types, we find evidence that generated explanations take advantage of models’ privileged access to their own internal mechanisms. Quantifying how explainer model capacity, target model complexity, task difficulty, etc. influence the emergence or manifestation of privileged access remains an important question for future work.
@@
In a general sense, our results suggest that objectives that enforce self-consistency (Anonymous Authors, 2026) between models’ behavior and explanations may provide an avenue towards more scalable interpretability methods, faithful explanations, and alignment guarantees. Generalized versions of our objectives could be applied to any method for generating ground-truth descriptions of model computation or behavior, e.g. to describe inferred user traits, surface adversarial triggers, or summarize influential training examples. These objectives could also be applied in the reverse direction, fine-tuning models so that their behavior matches generated explanations, which might be normatively correct even if inaccurate as self-descriptions. This could provide a new mechanism for “unsupervised alignment” without relying on large preference datasets or reinforcement learning procedures.
Impact Statement
This work introduces a way to train models to produce accurate self-descriptions of their internal procedures. In general, we anticipate that this ability can be used as a way for debugging and red-teaming models, understanding model decisions, and improving explanation accessibilty for non-expert users. We also showed that this ability enables a scalable path towards procuring feature descriptions in Section˜4; we anticipate that many other forms of self-descriptions are also significantly more data efficient than their analogous interpretability techniques. This can help us save on the compute and human costs of running these interpretability experiments, while enabling more people to have access to explanations. By democratizing model explanations, we enable greater trust and interactivity of future AI systems.
On the other hand, introspective capabilities paired with misaligned models and greater user trust could potentially pave the way for negative outcomes. First, more faithful self-verbalizations may encourage the building of oversight mechanisms that are over-reliant on these verbalizations, or mislead users into overestimating model trustworthiness. We believe self-verbalization will always remain complementary to interpretability techniques that operate on model internals or conduct explicit counterfactual experiments on models; these verbalizations give fast access into models internals, but must always be validated against more rigorous techniques in high-stakes scenarios. Second, and more philosophically, AI systems can gain a better understanding of how they work could potentially have greater deceptive capabilities. For example, AI systems may have a greater understanding of their training pipelines and the situationally context that they’re in (e.g. training/evaluation/deployment), allowing them to operate in one way during evaluation and another way during deployment (Laine et al., 2024).
Appendix A Feature Description: LM Judge Results
For SAE features, gold descriptions are available via Neuronpedia, so we use an LM judge to assess the similarity of the predicted description to the gold description, on a scale of 0 (no similarity) to 1 (very similar), in increments of 0.25. We use GPT-4.1-mini as the judge.
A.1 Prompt
@@ A.2 Human Validation
A.3 Error Analysis
We also perform a human qualitative analysis on 100 test set outputs from the explainer model that received a score of 0 from the LM judge in Table˜1. Our analysis reveals that sources of error extend beyond the explainer model itself and can be attributed to multiple stages of the evaluation pipeline, including (1) the original feature labels generated by the LM based on activation exemplars and (2) the validity of the LM judge’s assessment when comparing the explainers’ output against the original labels. Errors from these sources do not reflect failures of the explainer model.
Appendix B Feature Descriptions: Additional Experimental Details
B.1 Explainer models
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
B.2 Layerwise Nearest Neighbors
We also evaluate a nearest-neighbors baseline where, instead of retrieving from all training features from all layers, we only retrieve from the training features from the corresponding test layer. Let $S_{\text{train},\ell}$ denote training features from layer $\ell$. Our layerwise nearest neighbor baseline is thus:
$\mathcal{E}_{\text{NN-layer}}(v,\ell)=D\!\left(\arg\max_{v_{i}\in S_{\text{train},\ell}}\langle v_{i},v\rangle\right)$
where $D$ is the function mapping training features to their explanations.
Appendix C Feature Descriptions: Additional Results
C.1 Gemma-2-9B as Target Model
We replicate a subset of the evaluations in Table˜1 using Gemma-2-9B as the target model. We select Gemma-2-9B because its SAE features are also publicly available via GemmaScope (Lieberum et al., 2024), and downloadable via Neuronpedia (Lin, 2023). Results can be found in Table˜3.
@@ C.1 Gemma-2-9B as Target Model
C.2 Quantifying privileged access via activation alignment
@@
For each similarity metric and each explainer model (fixing 3.1-8B as the target model), we provide raw similarity values in Table˜4. We then plot explainer similarities vs. explainer performance for each of the four (metric, feature type) pairings, as described in Section˜3.2. Plots can be found in Figure˜7. We find that explainer performance tracks the activation similarity of the explainer, indicating alignment between activations predicts verbalization capability.
@@
C.3 Is Layer Annotation Necessary to Describe SAE Features?
@@
Appendix D Activation Patching: Additional Results
Target Explainer Exact Match Has-Changed F1 Content Match
Qwen3-8B Qwen3-8B 64.0±0.4 80.2±0.3 71.0±0.4
– activation 59.9±0.4 76.5±0.4 68.6±0.4
– layer 65.0±0.4 80.0±0.3 72.7±0.4
– token 63.9±0.4 79.4±0.3 71.7±0.4
Llama-3.1-8B 54.1±0.4 73.9±0.4 65.4±0.4
Qwen-3-8B (Untrained) 5.02±0.2 24.4±0.8 18.78±0.3
Llama-3.1-8B Llama-3.1-8B 48.6±0.7 77.5±0.6 54.6±0.7
– activation 45.2±0.7 73.9±0.6 50.8±0.7
– layer 49.7±0.7 76.9±0.6 55.7±0.7
– token 47.3±0.7 76.1±0.6 53.4±0.7
Qwen3-8B 41.7±0.7 75.1±0.6 47.6±0.7
Llama-3.1-8B (Untrained) 3.2±0.1 31.4±0.4 7.0±0.2
Table 5: Activation patching outcomes for all target models. We train Qwen3-8B and Llama-3.1-8B to verbalize their own activation patching outcomes, and compare against training the other model to explain their outcomes, and also against untrained versions of themselves. Scores (mean $\pm$ standard error) are reported for three metrics—exact match, has-changed F1, and content prediction. For each target model and metric, bold indicates no significant difference from the best entry (paired $t$-test, $p\geq 0.05$). We also conduct experiments ablating the activation, layer, and token information from the explainer’s input. Ablating layer and token each have neglible effect, potentially because they can already be inferred from the activation.
D.1 Llama-3.1-8B as Target Model
We report full activation patching results, including results of Llama-3.1-8B as the target model in Table˜5. These results complement Table˜2 in the main paper.
We observe consistent privileged access: Llama-3.1-8B is best at explaining its own activation patching outcomes, outperforming Qwen3-8B and the untrained baseline. Similarly, Qwen3-8B is best at explaining its own activation patching outcomes, outperforming Llama-3.1-8B and the untrained baseline.
D.2 Ablation Analysis
In addition to training each explainer model on each target model, we perform a set of ablations to isolate which input component(s) the explainer learned to condition on. The explainer input consists of three components: the activation $v$ to insert (– activations), the layers $\ell_{1:i}$ (– layer), and the token $t$ of the target model to patch into (– token). We train the explainer model with each aspect of the input ablated (conditioning only on the other two components) to generate the ground-truth explanations. [^13] Ablated prompts can be found in Section H.3. Results can be found in Table˜5. Our ablation studies suggest found that performance is generally robust to ablating the layer and token annotations in the input prompt. We believe this is because layer and token information is generally recoverable from the activation and input context alone.
As evidence for this, we train Llama-3.1-8B to decode the layer and token information from the input context and activation value alone, i.e. given the prompt
@@ D.2 Ablation Analysis
Appendix E Input Ablation: Additional Results
E.1 Full Results by Explainer
We report full input ablation results, including results of Llama-3.1-8B as a target model, grouped by explainer model rather than target model (unlike previous tables). This organization highlights the relative strength of different explainer models. These results complement Table˜2 in the main paper.
Target Explainer Exact Match Has-Changed F1 Content Match
Llama-3.1-8B Llama-3.1-8B 63.8±1.3 74.2±1.1 75.3±1.2
Qwen3-8B 58.1±1.3 70.5±0.0 71.6±1.2
Llama-3.1-8B Llama-3.1-8B (Untrained) 8.1±0.8 33.4±0.1 15.5±1.0
Qwen3-8B Qwen3-8B 83.4±1.0 87.0±1.1 90.6±0.8
Llama-3.1-8B 56.7±1.3 71.3±0.0 67.8±1.3
Qwen3-8B Qwen3-8B (Untrained) 8.9±1.5 44.4±1.6 35.3±2.5
Table 6: Input ablation results grouped by explainer model. We train Qwen3-8B and Llama-3.1-8B to verbalize their own input ablation outcomes, and compare against training to verbalize each other’s outcomes and untrained models. Scores (mean $\pm$ standard error) are reported for three metrics—exact match, has-changed F1, and content prediction. For each explainer model and metric, we bold all entries not significantly different from the best entry (paired $t$-test, $p\geq 0.05$). We find evidence for privileged access even though Llama-3.1-8B is a weaker explainer than Qwen3-8B overall.
Appendix F Explainer Model Prompts